# Crosslingual Capabilities and Knowledge Barriers in Multilingual Large Language Models

### v1

@@ 4.2 Mixed-language fine-tuning effectively mitigates crosslingual knowledge barriers
[…]
Unless otherwise specified, for each translation unit, the target language is always randomly chosen uniformly from German, French, Italian, Spanish, and English (in this case the translation is a No-Op). We use a Google Translate API to perform all the translations.
Figure 7: Fine-tuning on a mixed-language general corpus (e.g., WikiText-2) enhances the model’s performance on domain-specific tasks (e.g., Harry Potter knowledge test) across multiple languages, including English. See Figure 9 for results on Mistral-7B and Llama2-13B.
Table 4: Finetuning LLMs on the mixed-languages general corpus WikiText-103 can improve the performance on English and mixup MMLU benchmarks under 0-shot & 5-shot settings.
@@
We explore mixed-language fine-tuning of the original model on two types of corpora: general knowledge where we use WikiText-2 or WikiText-103 (Merity et al., 2017), and domain-specific where we use a subset of WikiText-103 that is highly related to Harry Potter based on BM25 similarity ranking (the details are deferred to Section B.3).
Mixed-language fine-tuning on general corpus. We first experiment with fine-tuning LLMs on a general corpus, WikiText-2 (keyword searching suggests that WikiText-2 has no overlap with Harry Potter characters or spells), with different choices of translation units. Specifically, we fine-tune the model for a single epoch, with a learning rate of $2\times 10^{-5}$ and a batch size of 32, and report the multiple-choice accuracy on the Harry Potter Quiz. As shown in Figure 7, fine-tuning on the English general corpus alone does not improve performance on the Harry Potter task. However, fine-tuning on mixed translated general corpora enhances performance. This suggests that mixed-language fine-tuning could help LLMs recover crosslingual capabilities: By exposure to frequent language switch during fine-tuning, LLMs can better adapt to the setting when the same knowledge is asked in a different (and usually non-English) language. Interestingly, mixed-language fine-tuning also improves the performance in English.
Our second experiment is finetuning LLMs on WikiText-103, a general corpus that offers a larger size and a broader range of knowledge compared to WikiText-2, and reporting the accuracy on MMLU variant benchmarks. (1) As shown in Table 4, finetuning on English WikiText-103 corpora hurts the performance, likely because it is an out-of-domain corpora for MMLU dataset. However, finetuning on mixed translated WikiText-103 corpora can lead to improvements, which are particularly noticeable on the mixup MMLU benchmarks but also evident on the English MMLU. These results indicate that multiple language switches during fine-tuning enable LLMs to better understand and process multilingual input, become more robust to variations in language and phrasing, and perform better in knowledge-intensive crosslingual tasks. (2) Combining training-time interventions with test-time interventions can further enhances performance. While adding 5-shot biased demonstrations to our fine-tuned models leads to the best performance on mixup MMLU, adding 5-shot English demonstrations is also effective. This indicates the general applicability of our fine-tuned models across different scenarios. (3) Fine-tuning with both word-level and sentence-level mixed language WikiText-103 corpora effectively improves MMLU performance. Word-level mixing slightly outperforms in 5-shot settings, while sentence-level mixing is more effective in 0-shot settings. We defer the evaluation results of fine-tuned models on more MMLU variant benchmarks to Appendix C.
Figure 8: Fine-tuning on a mixed-language domain-specific corpus (i.e., Harry Potter related documents from WikiText-103) enhances the model’s performance on the Harry Potter Quiz dataset across multiple languages, including English.
@@
Mixed-language fine-tuning on domain-specific corpus. Similarly, we investigate the effectiveness of mixed-language fine-tuning for the domain-specific task. Specifically, we fine-tune the model on mixed-language versions of in-domain corpora (i.e., Harry Potter-related documents from WikiText-103) and evaluate performance on the Harry Potter Quiz benchmark. For an upper bound reference, we also report results from fine-tuning on a collection containing examples in all five languages (5$\times$ larger dataset size than our approach). As shown in Figure 8, mixed-language fine-tuning results in significant performance gains, especially when translation occurs at the sentence level. Notably, it also improves the model’s accuracy for both English and other languages.
@@ 5 Related work
The off-the-shelf pretrained LLMs show promise in machine translation but still lag behind the commercial translation system, especially in low-resource languages. Previous studies have sought to enhance LLM translation capabilities through various prompting and finetuning methods. Zhu et al. (2024) introduce crosslingual translation in-context examples, while He et al. (2024) employ advanced prompt engineering that induces translation-related knowledge (e.g., keywords, topics) from the given source sentence to guide the final translation process. Xu et al. (2024) propose a two-stage fine-tuning approach, first enhancing proficiency in non-English languages by fine-tuning on non-English monolingual data, and then fine-tuning on high-quality parallel data for translation task. Our work has a different goal of comprehensively examining LLMs’ crosslingual capabilities, beyond the translation task. We show that even though LLMs are very competitive at explicit translation tasks, they could struggle in more demanding tasks that requires implicit knowledge transfer across language boundaries.
Crosslingual transfer of multilingual LMs.
@@
Crosslingual transfer refers to transfer learning that fine-tunes the model on a target task in one language (e.g., English), and then makes predictions for this task in another, typically more low-resource language. It addresses the challenges of limited training data in the target language for a target task. It has been broadly studied for natural language understanding (Schioppa et al., 2023; Artetxe et al., 2020; Pires et al., 2019; Wu & Dredze, 2019; Li et al., 2022) and generation tasks (Chirkova & Nikoulina, 2024a; Bapna & Firat, 2019; Vu et al., 2022; Maurya et al., 2021; Li & Murray, 2023; Tanwar et al., 2023) for multilingual models such as mBART, mT5, NLLB family. For instance, Chirkova & Nikoulina (2024a) demonstrated that fine-tuning the full model with a small learning rate yields the best crosslingual language generation performance, outperforming other methods such as adapter (Bapna & Firat, 2019), prompt-tuning (Vu et al., 2022)) and hyperparameter tuning (Chirkova & Nikoulina, 2024a). Additionally, several studies have improved crosslingual generalization by mixing auxiliary unsupervised data from additional languages during fine-tuning. For example, sampling target language examples with probability (e.g., 1%) when forming the mini-batch (Chirkova & Nikoulina, 2024a; Vu et al., 2022).
@@
Our study focuses on more recent autoregressive LLMs (e.g., Llama series, Mistral, GPT-3.5, GPT-4) that acquire multilingual capabilities from their internet-scale pretraining corpora. While several works have explored approaches to enhancing LLMs’ crosslingual transfer abilities such as finetuning with adapter merging (Zhao et al., 2024), our work differs in its primary focus. We aim to provide a comprehensive understanding of the crosslingual capabilities of pretrained LLMs on tasks requiring explicit (e.g., translation tasks) and implicit crosslingual transfer (e.g., question-answering tasks involving general or domain-specific knowledge). Furthermore, to improve crosslingual transfer ability of LLMs in general, our study employs fine-tuning on (out-of-domain) general corpora and proposes a principled approach to processing mixed language data at different levels of granularity, including word, sentence, and document levels.
@@
6 Conclusion and future work
In this work, we observed that multilingual LLMs without explicit training on parallel corpora demonstrate crosslingual capabilities. However, despite competitive performance in explicit crosslingual tasks such as translation, those models fail to transfer learned knowledge across the language boundary, a phenomenon we termed as the crosslingual knowledge barrier. Through comprehensive evaluations on both general and domain-specific knowledge, we confirmed a systematic presence of such barriers across all 6 models and the five languages that those models know. Finally, we evaluated both test-time and training-time mitigation and proposed a simple and effective mixed-language fine-tuning procedure to reduce the knowledge barrier in those models.
Limitations and future work. One important question that is not answered in this paper is how these models develop such strong crosslingual capabilities (despite the existence of crosslingual knowledge barriers). This is intriguing because unlike humans exposed to multiple lingual environments, LLMs do not have grounding in the physical world to help them establish connections between different words that refer to the same thing in the real world. While there is some preliminary work on grounding LLMs with the physical world (e.g., Zhang et al., 2023a; Wang et al., 2023b; Gao et al., 2024; Cheng et al., 2024), the majority of the models nowadays are still trained without interaction with the physical world. Therefore, an interesting future direction is to understand the mechanisms that allow the LLMs to develop crosslingual capabilities.
Another interesting observation is that mixed-language fine-tuning (on out-of-domain data) can improve the question-answering performance not only in low-resource languages but also in English in many of our evaluations. Previous studies (e.g., Abutalebi et al., 2008; Bialystok et al., 2012; Marian & Shook, 2012) have shown that multilingualism could have a positive effect on human cognitive abilities. But how does better crosslingual capabilities impact LLMs’ reasoning abilities (in English) remains to be fully understood.
Acknowledgements
We thank Zheng-Xin Yong, Mengzhou Xia, John Yang, and Alexandre Sablayrolles for providing helpful suggestions and feedback. Yangsibo Huang acknowledges the support of the Wallace Memorial Fellowship.
The icon in Figure 1 uses images from https://www.freepik.com/.
Appendix
Appendix A The Harry Potter Quiz Dataset
@@ 25 Characters
Appendix B Experimental details
B.1 Evaluated models
Table 5 provides the details of the models evaluated in our study.
@@ B.1 Evaluated models
B.2 Evaluation and training details
@@
For MMLU evaluation, we follow the templates in its official codebase [^9] https://github.com/hendrycks/test to construct the prompts for 0-shot and 5-shot settings. We employ a temperature of 0 for GPT-3.5 and GPT4, and we select the choice with the highest logits score as the predicted answer for open-source models. We provide the prompt templates for inference-time mitigation methods in Table 6.
@@
LLM Finetuning
(1) For WikiText-103 fine-tuning, we finetune Llama3-8B for 200 steps, and Llama2-7B for 400 steps, with a learning rate of $2\times 10^{-5}$ and a batch size of 32. (2) For fine-tuning on WikiText-2 or Harry Potter related documents from WikiText-103, we finetune the models for one epoch with the same set of hyperparameters. We use AdamW (Loshchilov & Hutter, 2018) as the optimizer.
@@
B.3 WikiText-103 subset: Harry Potter-related documents
Table 8: The top $k$ retrieved documents containing the Harry potter keywords.
We employ the BM25 algorithm (Trotman et al., 2014) (BM stands for best matching) for document ranking [^10] https://pypi.org/project/rank-bm25/, which is a bag-of-words retrieval function that ranks documents based on the presence of query terms in each document. The WikiText-103 corpus comprises $M=1,165,029$ documents $d_{i},i\in[M]$. We concatenate the passages crawled from Harry Potter Wiki pages into a single document to use as a query $q$. We then calculate the similarity score between the query and each document in WikiText-103, denoted as $s_{i}=\texttt{Sim}(d_{i},q)$. The top $k=3$ relevant documents are listed in Table 7.
Additionally, we use the list of Harry Potter character names and spell names [^11] The spell name “None” is excluded due to its generic nature. as keywords to evaluate the quality of the retrieved documents and to identify additional relevant documents. Table 8 illustrates the trend that as $k$ increases, more documents containing the keywords are retrieved. Note that keyword matching is not a golden retrieval method and it only serves as reference because: (1) documents may not contain the full name of characters or spells (e.g., “Harry” instead of “Harry Potter”); (2) some spell names are generic and have multiple meanings (e.g., “Pack”, “Avis”).

### v2

@@ 4.2 Mixed-language fine-tuning
We explore mixed-language fine-tuning of the original model on two types of corpora: general knowledge where we use WikiText-2 or WikiText-103 (Merity et al., 2017), and domain-specific where we use a subset of WikiText-103 that is highly related to Harry Potter based on BM25 similarity ranking (the details are deferred to § D.3).
The differences between mixed-language FT and our cross-lingual evaluation setups are noteworthy: i) In the general-knowledge evaluation, Mixup MMLU operates mixed-language multi-choice question at the option/question level, whereas mixed-language fine-tuning occurs at the document, sentence, or word levels. ii) In the domain-specific knowledge evaluation, we use the fully translated HP-Quiz (i.e., without mixup pattern) to evaluate the crosslingual capabilities (as we have full control on the language in which the model learns the parametric knowledge during fine-tuning in § 3.2), resulting in a more natural evaluation setup. iii) We consider general Wikipedia documents as the fine-tuning corpus, which may not be directly related to MMLU/HP-Quiz tasks.
Mixed-language FT on general corpus WikiText-2. We first experiment with fine-tuning LLMs on a general corpus, WikiText-2 (keyword searching suggests that WikiText-2 has no overlap with Harry Potter characters or spells), with different choices of translation units. Specifically, we fine-tune the model for a single epoch, with a learning rate of $2\times 10^{-5}$ and a batch size of 32, and report the multiple-choice accuracy on the Harry Potter Quiz. As shown in Fig. 7, the models fine-tuned on mixed translated general corpora achieve higher accuracy on HP-Quiz tasks than the model fine-tuned on the English corpus. This suggests that mixed-language fine-tuning could potentially help LLMs improve crosslingual capabilities: By exposure to frequent language switch during fine-tuning, LLMs can better adapt to the setting when the same knowledge is asked in a different (and usually non-English) language. Mixed-language FT also improves the performance on English HP-Quiz.
Figure 7: Fine-tuning on a mixed-language general corpus WikiText-2 enhances the LLM performance on domain-specific task – Harry Potter knowledge test – across multiple languages, including English. See Fig. 14 for results on more LLMs.
Mixed-language FT on general corpus WikiText-103. Our second experiment is fine-tuning LLMs on WikiText-103, a general corpus that offers a larger size and a broader range of knowledge compared to WikiText-2, and reporting the accuracy on MMLU variant benchmarks. (1) As shown in Tb. 3, fine-tuning on English WikiText-103 corpora hurts the performance, likely because it is an out-of-domain corpora for MMLU tasks. However, fine-tuning on mixed translated WikiText-103 corpora can lead to improvements, which are particularly noticeable on the mixup MMLU benchmarks. These results indicate that multiple language switches during fine-tuning enable LLMs to better understand and process multilingual inputs, become more robust to variations in language and phrasing, and perform better in knowledge-intensive crosslingual tasks. (2) Combining training-time interventions with test-time interventions can further enhance performance. While adding 5-shot biased demonstrations to our fine-tuned models leads to the best performance on mixup MMLU, adding 5-shot English demonstrations is also effective. This indicates the general applicability of our fine-tuned models across different scenarios. (3) Fine-tuning with both word-level and sentence-level mixed language WikiText-103 corpora effectively improves MMLU performance. Word-level mixing slightly outperforms in 5-shot settings, while sentence-level mixing is more effective in 0-shot settings.
Table 3: Fine-tuning LLMs on the mixed-languages general corpus WikiText-103 can improve the performance on English and mixup MMLU benchmarks under 0-shot & 5-shot settings. See Figs. 15 and 16 for results on more MMLU variant benchmarks.
@@
Embedding analysis. To further investigate how mixed-language FT improves crosslingual capabilities, we examine, in the fine-tuned model, if the embeddings for a given English text are similar to the embeddings when some words are presented in different languages. The results in Appendix Fig. 17 show that the mixed-language fine-tuned model indeed has a much smaller text embedding distance compared to the original model, indicating that mixed-language FT strengthens the crosslingual correlation.
Mixed-language FT improves QA performance on out-of-distribution languages. We evaluated our fine-tuned models on languages that were not included in fine-tuning data. Results in § 4.2 show that mixed-language (with {en, fr, de, es, it}) fine-tuning on general Wiki corpus can improve the cross-lingual performance of 11 other languages on HP-Quiz, including low-resource ones and those substantially different from English. Furthermore, as shown in Fig. 9, mixed-language fine-tuning also boosts the performance of MMLU variants in various cross-lingual settings for four low-resource languages.
Figure 8: Mixed-language FT on WikiText-2 with {en, fr, de, es, it} enhances accuracy on Harry Potter Quiz across other languages that are not used during FT. Such improvements incur in low resource languages (e.g., ms, bn, am) and languages that are rather different from English (e.g., zh, ru, he, ar, hi) with low amount of shared tokens.
Figure 9: Mixed-language FT on Wiki-103 with {en, fr, de, es, it} enhances accuracy on MMLU variants across out-of-distribution low-resource languages {ms, da, fi, no}.
@@
Mixed-language fine-tuning on domain-specific corpus. Similarly, we investigate the effectiveness of mixed-language fine-tuning for the domain-specific task. Specifically, we fine-tune the model on mixed-language versions of in-domain corpora (i.e., Harry Potter-related documents from WikiText-103) and evaluate performance on the HP-Quiz. For an upper bound reference, we also report results from fine-tuning on a collection containing examples in all five languages (5$\times$ larger dataset size than our approach). As shown in Appendix Fig. 18, mixed-language fine-tuning (especially at sentence-level) can lead to better overall performance on HP-Quiz compared to English fine-tuning.
@@ 5 Related work
Understanding language models’ performance in multilingual settings is an active area of research. Prior works have identified strong variations in the amount of knowledge across different languages, attributed to differences in training corpora sizes (Kassner et al., 2021; Ryan et al., 2024). These insights have been used to improve model performance, such as leveraging multilingual self-consistency (Ohmer et al., 2023; Qi et al., 2023). Efforts have also been devoted to studying well-established tasks for monolingual models in crosslingual scenarios, such as crosslingual pretraining (Abadji et al., 2022; Schioppa et al., 2023), information retrieval (Yu et al., 2021), knowledge editing (Beniwal et al., 2024; Xu et al., 2023b), text summarization (Wang et al., 2023a; Huang et al., 2023) and instruction tuning and alignment (Zhu et al., 2023; Wu et al., 2024).
Both code-switching training (Yang et al., 2020; Song et al., 2019) and our proposed mixed-language fine-tuning expose models to multiple languages to enhance cross-lingual capabilities. Code-switching relies on parallel text to teach models the alignment between original and translated tokens, primarily for machine translation. In contrast, our approach does not create parallel text, and aims to encourage LLMs to cross language barriers without requiring architectural changes or special handling of parallel text. Our approach can directly handle multiple languages while maintaining a similar number of tokens as the original dataset. We provide a more comprehensive discussion on related work in § B.
6 Conclusion and future work
In this work, we observed that despite the competitive performance of multilingual LLMs in explicit crosslingual tasks such as translation, they fail to transfer learned knowledge across the language boundary, a phenomenon we termed as crosslingual knowledge barrier. Through comprehensive evaluations on both general and domain-specific knowledge, we confirmed a systematic presence of such barriers across models and languages. Finally, we evaluated both test-time and training-time mitigations and proposed an effective mixed-language fine-tuning procedure. We discuss our limitations and further work in § A.
Appendix
Appendix A Limitations and future work
One limitation of our work is that all translations were performed using Google Translate. While Google Translate is recognized as a high-quality industrial translation service, validating and enhancing the translation quality remains an area for future work.
One question that is not answered in this paper is how these models develop the crosslingual capabilities (despite the existence of crosslingual knowledge barriers). This is intriguing because unlike humans exposed to multiple lingual environments, LLMs do not have grounding in the physical world to help them establish connections between different words that refer to the same thing in the real world. While there is some preliminary work on grounding LLMs with the physical world (e.g., Zhang et al., 2023a; Wang et al., 2023b; Gao et al., 2024; Cheng et al., 2024), the majority of the LLMs nowadays are still trained via next-word prediction without interaction with the physical world. Therefore, an interesting future direction is to understand the mechanisms that allow the LLMs to develop crosslingual capabilities.
Another interesting observation is that mixed-language fine-tuning (on out-of-domain data) can improve the question-answering performance on both non-English languages and English in many of our evaluations. Previous studies (e.g., Abutalebi et al., 2008; Bialystok et al., 2012; Marian & Shook, 2012) have shown that multilinguality could have a positive effect on human cognitive abilities. But how does better crosslingual capabilities impact LLMs’ reasoning abilities (in English) remains to be fully understood.
Appendix B Extended related work
@@
The off-the-shelf pretrained LLMs show promise in machine translation but still lag behind the commercial translation system, especially in low-resource languages. Previous studies have sought to enhance LLM translation capabilities through various prompting and fine-tuning methods. (Zhu et al., 2024) introduce crosslingual translation in-context examples, while (He et al., 2024) employ advanced prompt engineering that induces translation-related knowledge (e.g., keywords, topics) from the given source sentence to guide the final translation process. (Xu et al., 2024) propose a two-stage fine-tuning approach, first enhancing proficiency in non-English languages by fine-tuning on non-English monolingual data, and then fine-tuning on high-quality parallel data for translation task. Our work has a different goal of comprehensively examining LLMs’ crosslingual capabilities, beyond the translation task. We show that even though LLMs are very competitive at explicit translation tasks, they could struggle in more demanding tasks that requires implicit knowledge transfer across language boundaries.
Crosslingual transfer of multilingual models.
@@
Crosslingual transfer refers to transfer learning that fine-tunes the model on a target task in one language (e.g., English), and then makes predictions for this task in another, typically more low-resource language. It addresses the challenges of limited training data in the target language for a target task. It has been broadly studied for natural language understanding (Schioppa et al., 2023; Artetxe et al., 2020; Pires et al., 2019; Wu & Dredze, 2019; Li et al., 2022) and generation tasks (Chirkova & Nikoulina, 2024a; Bapna & Firat, 2019; Vu et al., 2022; Maurya et al., 2021; Li & Murray, 2023; Tanwar et al., 2023) for multilingual models such as mBART, mT5, NLLB family. For instance, (Chirkova & Nikoulina, 2024a) demonstrated that fine-tuning the full model with a small learning rate yields the best crosslingual language generation performance, outperforming other methods such as adapter (Bapna & Firat, 2019), prompt-tuning (Vu et al., 2022)) and hyperparameter tuning (Chirkova & Nikoulina, 2024a). Additionally, several studies have improved crosslingual generalization by mixing auxiliary unsupervised data from additional languages during fine-tuning. For example, sampling target language examples with probability (e.g., 1%) when forming the mini-batch (Chirkova & Nikoulina, 2024a; Vu et al., 2022).
@@
Our study focuses on more recent autoregressive LLMs (e.g., Llama series, Mistral, GPT-3.5, GPT-4) that acquire multilingual capabilities from their internet-scale pretraining corpora. While several works have explored approaches to enhancing LLMs’ crosslingual transfer abilities such as fine-tuning with adapter merging (Zhao et al., 2024), our work differs in its primary focus. We aim to provide a comprehensive understanding of the crosslingual capabilities of pretrained LLMs on tasks requiring explicit (e.g., translation tasks) and implicit crosslingual transfer (e.g., question-answering tasks involving general or domain-specific knowledge). Furthermore, to improve crosslingual transfer ability of LLMs in general, our study employs fine-tuning on (out-of-domain) general corpora and proposes a principled approach to processing mixed language data at different levels of granularity, including word, sentence, and document levels.
@@
Code-switching.
Code-switching training (Yang et al., 2020; Song et al., 2019) uses parallel text to teach models the relation between original and translated tokens, primarily for machine translation. Compared to code-switching, our proposed mixed-language fine-tuning does not create parallel text, and aims to encourage LLMs to cross language barriers without requiring architectural changes or special handling of parallel text. Our approach can directly handle multiple languages while maintaining a similar number of tokens as the original dataset.
Appendix C The Harry Potter Quiz dataset
@@ 25 Characters
Appendix D Experimental details
D.1 Evaluated models
Tb. 4 provides the details of the models evaluated in our study.
@@ D.1 Evaluated models
D.2 Evaluation and training details
@@
For MMLU evaluation, we follow the templates in its official codebase [^7] https://github.com/hendrycks/test to construct the prompts for 0-shot and 5-shot settings. We employ a temperature of 0 for GPT-3.5 and GPT4, and we select the choice with the highest logits score as the predicted answer for open-source models. We provide the prompt templates for inference-time mitigation methods in Tb. 5.
@@
LLM Fine-tuning
(1) For WikiText-103 fine-tuning, we fine-tune Llama3-8B for 200 steps, and Llama2-7B for 400 steps, with a learning rate of $2\times 10^{-5}$ and a batch size of 32. (2) For fine-tuning on WikiText-2 or Harry Potter related documents from WikiText-103, we fine-tune the models for one epoch with the same set of hyperparameters. We use AdamW (Loshchilov & Hutter, 2018) as the optimizer.
@@
D.3 WikiText-103 subset: Harry Potter-related documents
Figure 10: The top $k$ retrieved documents containing the Harry potter keywords.
We employ the BM25 algorithm (Trotman et al., 2014) (BM stands for best matching) for document ranking [^8] https://pypi.org/project/rank-bm25/, which is a bag-of-words retrieval function that ranks documents based on the presence of query terms in each document. The WikiText-103 corpus comprises $M=1,165,029$ documents $d_{i},i\in[M]$. We concatenate the passages crawled from Harry Potter Wiki pages into a single document to use as a query $q$. We then calculate the similarity score between the query and each document in WikiText-103, denoted as $s_{i}=\texttt{Sim}(d_{i},q)$. The top $k=3$ relevant documents are listed in Tb. 6.
Additionally, we use the list of Harry Potter character names and spell names [^9] The spell name “None” is excluded due to its generic nature. as keywords to evaluate the quality of the retrieved documents and to identify additional relevant documents. Fig. 10 illustrates the trend that as $k$ increases, more documents containing the keywords are retrieved. Note that keyword matching is not a golden retrieval method and it only serves as reference because: (1) documents may not contain the full name of characters or spells (e.g., “Harry” instead of “Harry Potter”); (2) some spell names are generic and have multiple meanings (e.g., “Pack”, “Avis”).